OpenStudy (babyslapmafro):

What is the formula for finding the volume of a solid generated when using the cylindrical shells method revolving about the x-axis?

OpenStudy (babyslapmafro):

The formula for rotation about the y-axis is... \[\int\limits_{a}^{b}2\pi xf(x)dx\] do i just replace x's with y's when rotating about the x-axis? \[\int\limits_{a}^{b}2\pi yf(y)dy\]